### v3

# MemBuilder: Reinforcing LLMs for Long-Term Memory Construction via Attributed Dense Rewards

Abstract
Maintaining consistency in long-term dialogues remains a fundamental challenge for LLMs, as standard retrieval mechanisms often fail to capture the temporal evolution of historical states. While memory-augmented frameworks offer a structured alternative, current systems rely on static prompting of closed-source models or suffer from ineffective training paradigms with sparse rewards. We introduce MemBuilder, a reinforcement learning framework that trains models to orchestrate multi-dimensional memory construction with attributed dense rewards. MemBuilder addresses two key challenges: (1) Sparse Trajectory-Level Rewards: we employ synthetic session-level question generation to provide dense intermediate rewards across extended trajectories; and (2) Multi-Dimensional Memory Attribution: we introduce contribution-aware gradient weighting that scales policy updates based on each component’s downstream impact. Experimental results show that MemBuilder enables a 4B-parameter model to outperform state-of-the-art closed-source baselines, exhibiting strong generalization across long-term dialogue benchmarks. [^1] Code available at https://github.com/Zh1yuShen/MemBuilder
MemBuilder: Reinforcing LLMs for Long-Term Memory Construction via Attributed Dense Rewards
Zhiyu Shen1thanks: Equal contribution.thanks: Work done during internship at Tencent Inc., Ziming Wu2∗thanks: Corresponding author., Fuming Lai2, Shaobing Lian2, Yanghui Rao1 1School of Computer Science and Engineering, Sun Yat-sen University, Guangzhou, China 2Tencent Inc. shenzhy23@mail2.sysu.edu.cn, raoyangh@mail.sysu.edu.cn {jimmyzmwu, fuminglai, lokilian}@tencent.com
1 Introduction
Memory-augmented frameworks have emerged as a promising approach for maintaining consistency in long-term dialogues, which must track evolving contexts and historical states over extended timelines. While Retrieval-Augmented Generation (RAG) facilitates access to external knowledge, it treats retrieval units as independent, static chunks—failing to capture how information evolves or which historical facts have been superseded (Liu et al., 2024; Gao et al., 2024). Memory-augmented frameworks address this by decomposing information prior to storage: events receive independent timestamps and semantic concepts are structured into discrete units. This shifts the computational burden from processing entangled tokens during inference to retrieving precise, “pre-digested" fragments. Recent implementations like Mem0 Chhikara et al. (2025), MIRIX Wang and Chen (2025), and MemGPT Packer et al. (2023) exemplify this approach, constructing external memory that evolves with each interaction. However, these prompting-based systems face three key limitations: they operate in an “open loop” without feedback on whether the constructed memories benefit downstream tasks; their performance degrades sharply when substituting less capable models; and their reliance on expensive closed-source models for the token-intensive memory construction phase makes deployment cost prohibitive at scale. This raises a critical question: Can we instead train a model to perform memory construction through direct supervision?
Figure 1: Sparse trajectory-level rewards (top) vs. our attributed dense session-level rewards (bottom). Dense rewards provide learning signals at each session rather than only at trajectory end.
Current training-based approaches, such as Memory-R1 (Yan et al., 2025) and Mem-$\alpha$ (Wang et al., 2025), attempt to address this learning gap but face two critical technical bottlenecks: 1) Sparse Trajectory-Level Rewards: In long-term dialogues, a single reward given at the end of a multi-session trajectory is too sparse. The model cannot discern which session’s memory operations contributed to the final outcome, making gradient updates noisy and learning ineffective. 2) Multi-Dimensional Memory Attribution: While the previous method adopts a multi-dimensional memory, all components share a global reward, failing to distinguish between operations on different types of memories, regardless of actual downstream impact.
We present MemBuilder, a framework for training models to construct long-term memory with attributed dense rewards. Our architecture utilizes a multi-dimensional memory design, comprising Core, Episodic, Semantic and Procedural components, and trains a single, lightweight 4B model to manage all four components. Our approach introduces two key technical contributions to address the limitations of existing RL-based memory construction methods (Figure 1). First, we employ dense session-level rewards. Unlike traditional methods that assign a single reward after processing all sessions, we leverage synthetic session-level question generation to provide immediate feedback after each session’s memory operations. Second, we introduce contribution-aware gradient weighting to resolve multi-dimensional memory attribution. Since all four memory components share a common reward, their individual contributions remain ambiguous. Our mechanism addresses this by scaling gradient updates based on the downstream utility of the constructed memories; specifically, the gradient impact of a memory operation is proportional to the usage of its corresponding memory component during retrieval. We integrate these techniques into Attributed Dense Rewards Policy Optimization (ADRPO). Extensive experiments conducted on different benchmarks demonstrate the effectiveness of our proposed framework. Specifically, MemBuilder achieves 84.23% on LoCoMo, surpassing the baselines including Claude 4.5 Sonnet under the same setting. Notably, our model is trained exclusively on LongMemEval yet generalizes effectively to out-of-distribution benchmarks with different dialogue structures and question types.
2 Related Works
2.1 Long-Term Memory Management in Conversational Agents
Maintaining coherent and personalized interactions over extended dialogues remains a fundamental challenge for LLM-based agents (Zhong et al., 2024; Packer et al., 2023). Recent benchmarks such as LoCoMo (Maharana et al., 2024), LongMemEval (Wu et al., 2025), and PerLTQA (Du et al., 2024) evaluate long-term memory through multi-session QA, temporal reasoning, and evolving user profiles. Early approaches addressed context limitations through position encoding modifications (Chen et al., 2023; Peng et al., 2024) or dedicated long-context training (Tworkowski et al., 2023; Bai et al., 2024), but these incur high computational costs and struggle to capture the temporal dynamics in multi-session dialogues (Liu et al., 2024). While RAG offers better scalability (Gao et al., 2024), its chunk-level retrieval lacks the temporal and semantic organization needed for complex long-term reasoning (Liu et al., 2024), motivating memory-augmented solutions that decompose information into structured units.
2.2 Prompting-Based Memory Frameworks
Inspired by the cognitive science distinctions among episodic, semantic, and procedural memory (Tulving, 1972), recent frameworks construct structured external memory for LLM agents (Sumers et al., 2024; Laird, 2012). Representative prompting-based implementations include MemGPT with its operating system-like memory hierarchy (Packer et al., 2023), Mem0 for personalized memory extraction (Chhikara et al., 2025), and MIRIX for multi-dimensional organization (Wang and Chen, 2025), and SCM for self-controlled memory management (Wang et al., 2026). Recent work like MMS (Zhang et al., 2025b) and RMM (Tan et al., 2025) further incorporates cognitive principles into memory design. However, these prompting-based frameworks rely on expensive closed-source models and operate without feedback on downstream utility.
2.3 Training-Based Approaches for Memory Construction
Training-based methods can be categorized by memory form. Latent memory approaches encode information into hidden states: MEM1 consolidates memory through internal state mechanisms via RL (Zhou et al., 2025), while MemGen generates latent memory tokens within the reasoning stream (Zhang et al., 2025c), and LongMem (Wang et al., 2023) uses a decoupled architecture with a frozen backbone as memory encoder and an adaptive side-network as memory retriever. Though efficient, these implicit representations sacrifice interpretability and fine-grained controllability.
Explicit memory approaches train models to manage structured external stores. Memory-R1 employs RL with sparse trajectory-level rewards (Yan et al., 2025), but lacks learning signals for dense memory operations. Mem-$\alpha$ trains multi-dimensional memory construction, generalizing from 30k to 400k+ tokens (Wang et al., 2025), yet applies a global reward across all memory operations regardless of their downstream impact. While RLVR with GRPO (Shao et al., 2024) has become widely adopted (Zhang et al., 2025a), sparse rewards remain insufficient for long-term dialogues, motivating our ADRPO.
3 Methodology
3.1 Problem Formulation
We address the task of long-term dialogue question answering. Given a sequence of conversation sessions $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$ with associated timestamps $\{t_{1},t_{2},\ldots,t_{n}\}$, and a question $q$ posed at time $t_{q}$ where $t_{q}>t_{n}$, the goal is to generate an accurate answer based on information distributed across the entire conversation history. Since concatenating all sessions typically exceeds context limits, we introduce an external memory bank $\mathcal{M}$ that compresses and organizes historical information for selective retrieval at inference time.
3.2 Multi-Dimensional Memory Architecture
To effectively manage long-term dialogue information, we design a multi-dimensional memory system that decomposes conversations into four specialized memory types, each handled by a role-specific prompt to the same LLM (Figure 2).
Figure 2: Multi-Dimensional Memory Architecture. Four memory types (Core, Episodic, Semantic, Procedural) are constructed during the Build Phase and selectively retrieved during the Answer Phase.
Memory Structure.
Our memory bank $\mathcal{M}$ consists of four components:
  • Core Memory $\mathcal{M}^{\textrm{core}}$: A fixed-size block storing persistent user profile information including identity, preferences, and key relationships. This memory is always included in the context during question answering. Detailed prompt templates are provided in Appendix I.
  • Episodic Memory $\mathcal{M}^{\textrm{epi}}$: Time-stamped event records capturing what happened and when. Each entry follows the format “YYYY-MM-DD: Event summary | Details”, enabling temporal reasoning.
  • Semantic Memory $\mathcal{M}^{\textrm{sem}}$: Factual knowledge about entities in the user’s life, such as people, places, and user-specific concepts. Common knowledge is explicitly excluded to avoid redundancy.
  • Procedural Memory $\mathcal{M}^{\textrm{proc}}$: Step-by-step processes, routines, and workflows mentioned in conversations, such as the user’s morning routine or problem-solving approach.
Given a new conversation session, all four memory types are processed simultaneously, each extracting memories according to its specialized perspective. Core Memory is maintained as a fixed block with automatic compression when capacity is exceeded. The other three memory types are stored in a vector database and retrieved via semantic similarity during question answering.
Memory Operations.
Since Core Memory operates on a single text block while the other three manage independent entries (Section 3.2), their action spaces differ accordingly:
$\mathcal{A}^{\textrm{core}}$ $=\{\textsc{Append},\textsc{Replace},\textsc{Rewrite}\}$ (1)
$\mathcal{A}^{\textrm{epi}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Merge}\}$ (2)
$\mathcal{A}^{\textrm{sem}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Skip}\}$ (3)
$\mathcal{A}^{\textrm{proc}}$ $=\{\textsc{Add},\textsc{Update}\}$ (4)
At session $\tau$, given state $\mathit{State}_{\tau}=(\mathcal{M}_{\tau-1},s_{\tau})$, the LLM selects an action $a\in\mathcal{A}^{(m)}$ for each memory type and generates the corresponding memory content.
For Core Memory, Append adds new information to the block, Replace updates specific fragments, and Rewrite reorganizes the entire block. For the other three types, Add creates a new entry, and Skip bypasses common knowledge already captured in the model’s parameters.
Unlike prior memory systems that delete old entries and replace them with new ones (Chhikara et al., 2025; Wang and Chen, 2025), we introduce two operations that preserve temporal history. The Update operation creates a new entry with a fresh timestamp that explicitly references the previous entry, rather than overwriting it, enabling the model to trace how information evolved. The Merge operation synthesizes multiple related events into a conclusion spanning a time range while preserving references to the original events as evidence, pre-computing complex temporal reasoning to reduce the burden during question answering.
This architecture transforms unstructured dialogues into organized, queryable memory. The remaining challenge is how to train the LLM to construct memory that maximizes downstream QA performance.
3.3 Supervised Fine-Tuning
While multi-dimensional memory architectures can offer finer-grained control (Wang and Chen, 2025), lightweight models, such as Qwen3-4B employed in our framework, often struggle with direct multi-dimensional memory construction, frequently producing invalid actions—such as malformed JSON structures, missing required fields (e.g., episodic entries without timestamps), or misuse of undefined operation types—that impede effective RL exploration. To address this cold-start problem, we collect expert trajectories $\{(\mathit{State}_{t},a_{t})\}_{t=1}^{n}$ using Claude 4.5 Sonnet. This phase stabilizes the model’s output format, providing a viable baseline for subsequent training. However, SFT, as a form of imitation learning, trains the model to replicate expert memory operations but does not directly optimize for downstream QA accuracy; we further employ RL to maximize the model’s ultimate utility in QA contexts.
Figure 3: ADRPO training pipeline. Each session’s memory rollouts are evaluated via synthetic QA, with gradients weighted by each memory component’s downstream contribution.
3.4 Attributed Dense Rewards Policy Optimization (ADRPO)
While SFT enables valid action generation, the resulting policy lacks optimization for downstream QA utility. We introduce a reinforcement learning algorithm that addresses two key challenges: sparse trajectory-level rewards in long-term dialogues, and multi-dimensional memory attribution among memory components with varying downstream impacts. Figure 3 illustrates the ADRPO training pipeline.
3.4.1 Dense Session-Level Rewards via Synthetic Session-level QA
Prior RL approaches for memory construction (Yan et al., 2025; Wang et al., 2025) assign a single reward at the trajectory’s end based on the final QA result. For dialogues spanning dozens of sessions, this provides no learning signal for dense memory operations.
We address this through synthetic session-level QA that evaluates memory quality at each step. Before the RL training, for each session $\tau$, we retrieve the top-$k$ memories from $\mathcal{M}_{\tau-1}$ most similar to the session $s_{\tau}$, and let an expert model with $(s_{\tau},\mathcal{M}_{\textrm{retrieved}})$ generate $J$ question-answer pairs $\{(q_{j},\mathit{ans}_{j})\}_{j=1}^{J}$ targeting information in $s_{\tau}$ or its connections to retrieved memory.
During the RL training at $s_{\tau}$, we sample $N$ rollouts. Each rollout $i$ produces memory operations for all four memory types, yielding a candidate memory bank $\mathcal{M}_{\tau}^{(i)}$. A capable model answers pre-generated questions by retrieving from $\mathcal{M}_{\tau}^{(i)}$, and an LLM judge assesses correctness against ground-truth $\mathit{ans}_{j}$. The task reward measures memory construction quality as the average QA accuracy:
$r^{\text{task}}=\frac{1}{J}\sum_{j=1}^{J}\mathbbm{1}[\text{correct}(q_{j})]$ (5)
The final reward incorporates two regularization terms:
$r=\mathbbm{1}[\text{valid}]\cdot r^{\text{task}}\cdot(1-\lambda\cdot\ell)$ (6)
Format Validity.
The indicator $\mathbbm{1}[\text{valid}]$ acts as a gate: outputs with malformed JSON structure, missing required fields, or undefined actions receive zero reward regardless of content quality.
Length Penalty.
[…]
$\ell^{(\text{core})}=\begin{cases}0&\textrm{if }\Delta_{\text{core}}\leq\theta_{\min}\\[3.0pt] \frac{\Delta_{\text{core}}-\theta_{\min}}{\theta_{\max}-\theta_{\min}}&\textrm{if }\Delta_{\text{core}}\in(\theta_{\min},\theta_{\max})\\[3.0pt] 1&\textrm{if }\Delta_{\text{core}}\geq\theta_{\max}\end{cases}$ (7)
where $\theta_{\min}$ and $\theta_{\max}$ are the penalty-free and full-penalty thresholds. For the other memory types, let $\rho=|\mathcal{M}^{(m)}_{\text{new}}|/|\hat{\mathcal{M}}^{(m)}|$ and $\Delta=\big||\mathcal{M}^{(m)}_{\text{new}}|-|\hat{\mathcal{M}}^{(m)}|\big|$:
$\ell^{(m)}=\begin{cases}0&\textrm{if }\Delta<\delta\textrm{ or }\rho\in[\gamma_{l},\gamma_{u}]\\[3.0pt] \frac{\rho-\gamma_{u}}{\gamma_{\max}-\gamma_{u}}&\textrm{if }\rho\in(\gamma_{u},\gamma_{\max}]\\[3.0pt] \frac{\gamma_{l}-\rho}{\gamma_{l}-\gamma_{\min}}&\textrm{if }\rho\in[\gamma_{\min},\gamma_{l})\\[3.0pt] 1&\textrm{otherwise}\end{cases}$ (8)
where $\delta$ is the minimum difference threshold, $\gamma_{l},\gamma_{u}]$ the tolerance range, and $\gamma_{\min}$, $\gamma_{\max}$ the full-penalty boundaries.
3.4.2 Contribution-Aware Gradient Weighting
Within each rollout, actions operating on the four memory dimensions contribute to a shared memory bank and receive a global reward. However, the functional impact on downstream QA performance varies significantly across memory types; for instance, Episodic Memories may be frequently retrieved while Procedural Memories remain unused. To account for these discrepancies, we dynamically amplify gradient updates based on each component’s downstream utility. Since vanilla GRPO computes a single advantage per rollout and applies it uniformly to all components, it cannot distinguish their individual contributions—a challenge analogous to credit assignment in cooperative multi-agent settings (Foerster et al., 2018; Li et al., 2022). Drawing on prior work that scales gradient updates by downstream utility to resolve such attribution ambiguity (Li et al., 2022; Ren et al., 2018), we dynamically amplify gradient updates based on each component’s retrieval-based contribution.
During QA evaluation, we record retrieval counts $h^{(m)}$ for each memory type $m\in\{\textrm{epi},\textrm{sem},\textrm{proc}\}$ across all questions. The dominant contributing type is:
$d=\arg\max_{m\in\{\textrm{epi},\textrm{sem},\textrm{proc}\}}h^{(m)}$ (9)
Gradient weights are assigned as:
$w^{(m)}=\begin{cases}\alpha&\textrm{if }m=d\\ 1&\textrm{otherwise}\end{cases}$ (10)
where $\alpha>1$ amplifies updates for the dominant contributor. Core Memory, which is always included in the context rather than retrieved, receives a fixed weight $w^{(\textrm{core})}=1$. This mechanism ensures that memory types whose entries directly contributed to successful QA receive proportionally stronger reinforcement.
3.4.3 Training Objective
We formulate the ADRPO training objective by extending GRPO (Shao et al., 2024) with the attributed session-level reward. At each session $\tau$, we sample $N$ rollouts from the current policy. Each rollout $i$ invokes the model four times in parallel, producing memory operations $a_{i}=(a_{i}^{\text{core}},a_{i}^{\text{epi}},a_{i}^{\text{sem}},a_{i}^{\text{proc}})$ for each memory type. All four memory types share the session-level reward $r_{i}$ defined in Eq. 6, but receive differentiated gradient weights $w^{(m)}$ based on their retrieval-based attribution.
Advantages are computed via within-group normalization:
$A_{i}=\frac{r_{i}-\mu}{\sigma+\epsilon}=\frac{\mathbbm{1}[\text{valid}_{i}]\cdot r^{\text{task}}_{i}\cdot(1-\lambda\ell_{i})-\mu}{\sigma+\epsilon}$ (11)
where $\mu$ and $\sigma$ are computed over the $N$ rollouts, and $\epsilon$ is a small constant for numerical stability. The training objective is:
$\mathcal{J}(\theta)=\mathbb{E}\Bigg[\sum_{m}\frac{1}{|a_{i}^{(m)}|}\sum_{k=1}^{|a_{i}^{(m)}|}\min\Big(w^{(m)}\rho_{i,k}^{(m)}A_{i},\\ w^{(m)}\cdot\text{clip}(\rho_{i,k}^{(m)},1\!-\!\epsilon,1\!+\!\epsilon)A_{i}\Big)\Bigg]-\beta\cdot D_{\textrm{KL}}(\pi_{\theta}\|\pi_{\textrm{ref}})$ (12)
where $\rho_{i,k}^{(m)}=\pi_{\theta}/\pi_{\textrm{ref}}$ is the importance ratio for the $k$-th token of memory type $m$’s output. The contribution-aware weights $w^{(m)}$ scale both terms within the $\min$ operator, uniformly amplifying the objective for high-impact components while preserving the clipping mechanism for stability.
4 Experiments
4.1 Experimental Setup
Datasets.
We evaluate MemBuilder on three long-term dialogue benchmarks: LongMemEval (Wu et al., 2025), which consists of user-assistant chat histories designed to evaluate the long-term memory capabilities of chat assistants; LoCoMo (Maharana et al., 2024), which contains human-human conversations between fictional personas grounded on temporal event graphs, spanning up to 35 sessions; and PerLTQA (Du et al., 2024), a dataset featuring 141 characters with rich personal profiles, social relationships, and life events. We train exclusively on the LongMemEval subset. LoCoMo and PerLTQA serve as OOD test sets that differ in both dialogue format and domain. Detailed statistics and data construction procedures are provided in Appendix A.
Baselines.
We compare against: (1) RAG-based: Following common practice in RAG-based systems (Mastra, 2025), we implement two retrieval granularities: RAG-Session chunks dialogues at session boundaries and retrieves complete sessions, while RAG-Utterance embeds individual utterances for fine-grained matching but returns the containing session to preserve conversational context; (2) Memory frameworks: Mem0 (Chhikara et al., 2025) and MIRIX (Wang and Chen, 2025); (3) Training-based: Memory-R1 (Yan et al., 2025), whose results are taken from the original paper due to unavailable code.
Implementation Details.
We use Qwen3-4B-Instruct-2507 as our base model. GPT-4.1 serves as the LLM judge for evaluation. All retrieval uses text-embedding-3-small. To isolate memory construction quality, we fix the answer model to Claude 4.5 Sonnet unless otherwise specified. Training hyperparameters are provided in Appendix B and baseline details in Appendix D. Detailed configuration including embedding settings and action formats is in Appendix C. A detailed cost breakdown is provided in Appendix F.
4.2 Training Data Construction
We use LongMemEval as our sole training source, sampling 50 dialogues for SFT and a separate 50 for RL. All other benchmarks serve as OOD test sets.
SFT Dataset.
The 50 SFT dialogues comprise approximately 2,400 sessions. We collect expert trajectories using Claude 4.5 Sonnet by processing sessions sequentially: at step $k$, the model receives the new session $s_{k}$ along with the top-20 memories retrieved from the step k1k{-}1 memory bank, then generates memory operations for all four types. After each step, newly created memories are embedded and added to the vector database. Since each session produces four memory type outputs trained as separate examples, this yields 9,600 training samples with an average input length of 6,000 tokens and output length of 780 tokens.
RL Dataset.
The 50 RL dialogues similarly comprise approximately 2,400 sessions. For each session $\tau$, we construct the policy model’s input by retrieving the top-20 most relevant memories from the step $\tau{-}1$ vector database, concatenated with the new session $s_{\tau}$, mirroring the inference-time setup.
To enable dense session-level rewards, we generate 5 QA pairs per session (12,000 pairs total) using Claude 4.5 Opus. The generation model receives the new session along with the retrieved memories, allowing it to create questions that test both current session retention and connections to prior history. Questions span three types: single-session (testing current session content), multi-session (requiring cross-session aggregation), and temporal-reasoning (involving time-based inference).
Reward Computation.
During GRPO training (8 rollouts per session, 5 epochs), each rollout $i$ at session $\tau$ produces candidate memory bank $M^{(i)}_{\tau}$. GPT-4.1-mini answers the pre-generated questions by retrieving from $M^{(i)}_{\tau}$, and an LLM judge (GPT-4.1-mini) evaluates correctness against ground-truth answers. The average accuracy across 5 questions yields the task reward $r_{\textrm{task}}$ (Eq. 6). Within each rollout, all four memory types share this reward, differentiated through contribution-aware gradient weighting (Section 3.4.2). This produces 96,000 session-rollouts for policy optimization. A human evaluation on 100 sampled QA pairs confirms that 92% of generated answers are factually correct (Appendix H.2).
4.3 Main Results
Table 1 presents performance across three benchmarks. To isolate the effect of memory construction quality, we set the answer model to Claude 4.5 Sonnet across all methods and compare three categories of approaches: retrieval-based methods, prompting-based frameworks, and training-based methods. Note that Memory-R1 trains Llama-3.1-8B-Instruct as both the memory construction and answer model. For a fairer comparison, we also evaluate our method with Qwen3-4B as the answer model (Table 2), which still achieves 82.00% on LoCoMo, significantly outperforming Memory-R1 (62.74%).
Method LoCoMo LongMemEval PerLTQA
Retrieval-based Methods
RAG-Session 70.35 66.75 79.21
RAG-Utterance 74.87 69.00 77.23
Prompting-based Memory Construction
Mem0 51.64 47.00 62.04
MIRIX 77.48 73.25 83.11
Ours (GPT-4.1) 79.91 78.50 91.74
Ours (QwQ-32B) 77.47 76.00 88.96
Ours (Claude 4.5 Sonnet) 82.61 85.50 92.59
Training-based Memory Construction
Memory-R1 62.74 - -
Ours (Qwen3-4B) 68.07 56.00 76.85
   + SFT 81.74 84.25 91.67
   + RL 79.31 62.75 82.19
   + SFT + RL 84.23 85.75 93.14
Results from the original paper with a different answer model.
Table 1: Performance comparison of different memory construction methods. “Ours” denotes our memory architecture with different memory construction models.
Our method achieves SOTA performance across all three benchmarks. On LoCoMo, our trained Qwen3-4B model achieves 84.23%, surpassing the best prompting-based framework MIRIX (77.48%) by 6.75 percentage points and outperforming Claude 4.5 Sonnet as the memory construction model (82.61%). Similar trends are observed on LongMemEval (85.75%) and PerLTQA (93.14%), where our method also outperforms all baselines including Claude 4.5 Sonnet. These results demonstrate that, although memory construction requires frequent model invocations across sessions, a well-trained 4B model can effectively replace expensive closed-source APIs. Notably, our model is trained exclusively on LongMemEval, yet achieves strong performance on LoCoMo and PerLTQA, demonstrating robust generalization to OOD benchmarks with different dialogue structures and question types.
The training stage ablation reveals the complementary roles of SFT and RL. SFT alone improves the base model from 68.07% to 81.74% by enabling valid multi-dimensional outputs, while RL further boosts performance to 84.23% by optimizing for downstream QA utility. Notably, RL without SFT (79.31%) underperforms SFT alone, confirming that supervised fine-tuning is essential to address the cold-start problem before effective RL exploration can proceed.
4.4 Ablation Studies
We conduct ablation experiments to analyze the contribution of our key design choices. Implementation details are provided in Appendix E.
4.4.1 Effect of Gradient Weighting
Figure 4: Training curves with different gradient weighting coefficients $\alpha\in\{1,2,4,8,16\}$ on LoCoMo.
To investigate the effect of contribution-aware gradient weighting (Section 3.4.2), we vary the weighting coefficient $\alpha$ that amplifies updates for the dominant contributing memory type. We conduct this ablation on a reduced training set for efficiency. As shown in Figure 4, performance improves as $\alpha$ increases from 1 (no weighting, 82.82%) to 4 (84.04%), confirming that attributing credit to high-contribution memory types enhances the final model performance. However, excessively large $\alpha$ values degrade performance due to gradient imbalance among memory types, with the optimal value at $\alpha=4$. A human evaluation further confirms that retrieval frequency is a reliable attribution proxy, with the dominant memory type aligning with the actual information source in 95% of cases (Appendix H.1).
4.4.2 Effect of Dense Rewards
Figure 5: Effect of reward density on LoCoMo accuracy. The x-axis indicates the fraction of sessions receiving task rewards during training.
To validate the importance of dense session-level rewards, we vary the reward density by providing task rewards to only a fraction of sessions during training. As shown in Figure 5, under the same number of training epochs, model performance degrades consistently as the reward becomes sparser. When reward density drops to 1/8, performance falls below the SFT baseline (80.92% vs 81.74%).
These results reveal that sparse rewards not only slow convergence but can also be worse than using SFT alone, explaining why prior sparse-reward approaches achieve limited gains despite employing larger base models.
4.4.3 Answer Model Generalization
Answer Model LoCoMo LongMemEval PerLTQA
Claude 4.5 Sonnet 84.23 85.75 93.14
GPT-4.1 81.51 82.50 91.83
Qwen3-4B Base 74.61 75.00 83.93
Qwen3-4B Ours 81.12 83.00 91.19
Table 2: Performance with different answer models using memory constructed by our trained Qwen3-4B model. “Qwen3-4B Base” denotes the base model, while “Qwen3-4B Ours” denotes our trained model.
To evaluate whether the constructed memory generalizes across different answer models, we fix the memory construction model to our trained Qwen3-4B and vary the answer model. Table 2 shows that our memory maintains high quality across answer models of varying capabilities. We further investigate the feasibility of replacing closed-source answer models with open-source alternatives in Appendix H.3, where results show that Qwen3-30B-A3B achieves performance comparable to GPT-4.1-mini on LoCoMo when evaluated on the same high-quality memory.
Interestingly, when using our trained Qwen3-4B as an answer model, accuracy improves from 74.61% to 81.12% over the base model and remains competitive with GPT-4.1, suggesting RL training creates implicit alignment between memory structure and the model’s reasoning patterns.
4.5 Further Analysis
4.5.1 Performance by Question Type
Method SingleHop MultiHop OpenDomain Temporal Adversarial
RAG-Utterance 68.75 51.35 84.29 69.23 85.11
Memory-R1 59.83 53.01 68.78 51.55 -
Ours 82.27 77.88 84.66 71.71 90.58
Table 3: Performance breakdown by question type on LoCoMo.
Table 3 details performance across LoCoMo question categories. Our method achieves the largest gains on MultiHop questions (77.88% vs 53.01% for Memory-R1, +24.87pp) and Temporal questions (71.71% vs 51.55%, +20.16pp), both of which require synthesizing information across multiple sessions. On Adversarial questions, our method achieves 90.58%, demonstrating robustness against misleading information.
4.5.2 Action Distribution Analysis
Figure 6: Action distribution across training stages (Base, SFT, RL) for four memory types.
Figure 6 visualizes how action distributions evolve across training stages. Concrete examples illustrating these behavioral changes are provided in Appendix G.
The most notable change is Core Memory’s shift from Append (98.1% $\rightarrow$ 12.9%) to Replace (13.3% $\rightarrow$ 87.1%). Our analysis of the generated outputs demonstrates that the model learns to perform targeted updates to specific fields rather than appending at the end. The RL training also teaches the model to be more selective at generation time rather than relying on post-hoc filtering. For Semantic Memory, Skip operations decrease, indicating that the model directly outputs relevant facts rather than enumerating candidates and then excluding. We also observe divergent Update behavior: usage increases for Episodic and Semantic Memories but drops to zero for Procedural Memory, suggesting that evolving information (e.g., events and facts) benefits from explicit update chains while procedural knowledge is better maintained by adding discrete new entries.
4.5.3 Training Dynamics
Figure 7: Training dynamics: (a) overall reward trend, (b) rewards by memory type, and (c) response length by memory type. All metrics show stable improvement without reward hacking.
Figure 7 illustrates the training dynamics of our RL process. (a) Both task reward (QA accuracy) and mean reward (with length penalty) improve steadily, indicating effective learning. (b) All four memory types show consistent reward growth. (c) Response lengths remain stable throughout training, confirming that the length penalty prevents reward hacking through verbose outputs.
5 Conclusion
We presented MemBuilder, a reinforcement learning framework for multi-dimensional memory construction in long-term dialogues. By introducing ADRPO, Qwen3-4B achieves 84.23% on LoCoMo, surpassing prompting-based frameworks using expensive closed-source models and generalizing effectively to OOD benchmarks. Our results demonstrate that memory construction can be handled by lightweight open-source models with appropriate training.
Limitations
To isolate the impact of memory construction quality and ensure fair comparison across different methods, our evaluation relies on a fixed closed-source model (Claude 4.5 Sonnet) for question answering, which incurs API costs during evaluation. However, as shown in Table 2, our constructed memory generalizes well across different answer models, suggesting that practitioners can substitute with capable open-source alternatives for cost-sensitive deployments. Furthermore, although we employ Claude 4.5 Opus for synthetic question generation, the generated QA pairs may still contain occasional inaccuracies or ambiguities. Despite this, our experimental results demonstrate that training with these synthetic questions substantially improves performance over sparse reward baselines.
Appendix A Dataset Details
A.1 Benchmark Datasets
We evaluate on three long-term dialogue benchmarks that differ in dialogue format, domain, and data organization.
LongMemEval consists of user-assistant chat histories designed to evaluate the long-term memory capabilities of chat assistants. The dataset contains 500 independent questions, each with its own dialogue context, averaging 40 sessions and approximately 115K tokens. Questions test five core memory abilities: information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention. This benchmark serves as both our training source and in-distribution evaluation set.
LoCoMo contains human-human conversations between fictional personas, grounded on temporal event graphs. The dataset includes 10 dialogues with an average of 27 sessions (range: 19–32) and 14K tokens per dialogue. Each dialogue is associated with multiple questions, totaling 1,986 questions across five types: SingleHop (282), MultiHop (321), OpenDomain (841), Temporal (96), and Adversarial (446). Unlike LongMemEval’s one-question-per-context format, LoCoMo tests memory systems on shared dialogue contexts with diverse question types.
PerLTQA is a dataset featuring 141 fictional characters, including 30 protagonists with rich personal profiles, social relationships, and life events. Questions are designed around the 30 protagonists, totaling 8,593 questions across five types: factual, reasoning, other, yes/no, and temporal. This dataset requires integrating Episodic Memories with Semantic Memories about characters.
Both LoCoMo and PerLTQA serve as out-of-distribution test sets, differing from LongMemEval in dialogue format (human-human vs. user-assistant), data organization (shared context vs. independent context), and domain.
Appendix B Training Details
B.1 SFT Training
We perform supervised fine-tuning using LlamaFactory (Zheng et al., 2024). The training data consists of expert trajectories where each example pairs a session input (retrieved memories + new session) with the expert model’s memory operations for all four memory types. We use Qwen3-4B-Instruct-2507 as the base model with learning rate $5\times 10^{-7}$, batch size 4, and train for 10 epochs.
B.2 RL Training
We implement ADRPO by extending the verl framework (Sheng et al., 2025) with contribution-aware gradient weighting and session-level reward computation. Starting from the SFT checkpoint, we train with a learning rate $1\times 10^{-6}$ and a batch size of 128. The rollout number is set to 8, and the clipping parameter $\epsilon$ is set to 0.2. For contribution-aware gradient weighting, we use $\alpha=4$ based on ablation results (Figure 4). The length penalty coefficient $\lambda$ is set to 0.8, with Core memory thresholds $\theta_{\min}=150$ and $\theta_{\max}=400$, and other memory parameters $\delta=200$ and $\gamma_{l},\gamma_{u}]=[0.5,1.3]$. Training runs for 5 epochs on 32 H20 GPUs (4 nodes), taking approximately 70 hours.
Appendix C Framework Configuration
C.1 Embedding and Retrieval
We use text-embedding-3-small as the embedding model for all retrieval operations throughout the framework. For memory construction, we retrieve the top-20 most relevant old memories to provide context for the current session. For QA answering, we similarly retrieve the top-10 memories from the final memory bank. The RAG baselines (RAG-Session and RAG-Utterance) retrieve the top-5 chunks. For synthetic QA generation during RL dataset construction, we retrieve the top-20 memories to provide richer historical context.
C.2 Memory Configuration
Core memory
is maintained as a single text block with a maximum capacity of 5,000 characters. When the content exceeds this limit after an Append or Replace operation, the same policy model is prompted to compress the content while preserving essential information, including user identity, key relationships, personality traits, important preferences, and long-term goals. The compression removes redundant descriptions, minor details, and verbose explanations. If the first compression attempt still exceeds the limit, a second, more aggressive compression pass is performed.
Episodic, Semantic, and Procedural Memories
are stored as individual entries in a vector database with no explicit size limit. Each entry is embedded independently for retrieval.
C.3 Agent Action Format
We refer to the role-specific prompt for each memory type as an agent. For each memory type, the model receives the current session along with relevant retrieved memories and outputs a JSON object specifying the action type and content. Below, we describe the action format for each agent.
Core memory Agent.
The Core memory Agent manages persistent user information, including identity, preferences, personality traits, and key relationships. It outputs one of three operations:
  • APPEND: Add new information to the existing Core memory block (used when capacity <<90%).
  • REPLACE: Update specific outdated or incorrect information by specifying old and new text.
  • REWRITE: Reorganize and consolidate the entire block (used when capacity >>90% or when major updates are needed).
Example output:
{"operation": "APPEND",
"content": "Works as a software engineer at Google, specializing in machine learning"}
Episodic Memory Agent.
The Episodic Memory Agent manages time-ordered event memories. Each entry includes a timestamp, summary, and detailed description capturing who, what, when, where, and why. It outputs operations from:
  • ADD: Create a new event entry not currently in memory.
  • UPDATE: Add a new related event that references previous events (old versions remain for history).
  • MERGE: Combine multiple related events into a timeline with a timestamp range, drawing conclusions from patterns (old versions remain for history).
Example output:
{"operations": [
{"action": "ADD",
"memory": "2024-03-15: Started new job at startup | Details: First day at TechCorp as senior engineer, met team lead Sarah..."}
Semantic Memory Agent.
The Semantic Memory Agent manages conceptual knowledge about people, places, objects, and concepts in the user’s life. It explicitly skips common knowledge already captured in the model’s parameters. Operations include:
  • ADD: Create an entry for a new concept, person, or object.
  • UPDATE: Add new information to an existing concept.
  • SKIP: Bypass common knowledge or already fully captured information.
Example output:
{"operations": [
{"action": "ADD",
"memory": "Sarah (colleague) - Career: Team lead at TechCorp, 5 years experience in ML..."},
{"action": "SKIP",
"reason": "Common knowledge about Python"}
Procedural Memory Agent.
The Procedural Memory Agent manages step-by-step processes, workflows, and instructions. Each entry includes a description, numbered steps, and optional context. Operations include:
  • ADD: Create a new procedure entry.
  • UPDATE: Modify an existing procedure with new information.
Example output:
{"operations": [
{"action": "ADD",
"memory": "Morning workout routine | Steps: 1. 10min stretching 2. 30min jogging 3. 15min core exercises | Context: Daily routine before work"}
Validity Criteria.
An action is considered valid if: (1) the JSON structure is well-formed, (2) the action type is defined for that agent, (3) all required fields are present, and (4) for UPDATE/MERGE operations, referenced entries exist in the current memory bank. Invalid actions receive zero reward regardless of content quality.
Appendix D Baseline Implementation
D.1 RAG Baselines
RAG-Session segments dialogues at session boundaries, treating each session as a retrieval unit. Given a question, we retrieve the top-5 most similar sessions using text-embedding-3-small and provide them as context to the answer model. RAG-Utterance segments dialogues at the utterance level, treating each user-assistant turn pair as a retrieval unit. Given a question, we retrieve the top-5 most similar utterances using text-embedding-3-small and provide them as context to the answer model.
D.2 Memory Frameworks
For fair comparison, we evaluate all memory frameworks using Claude 4.5 Sonnet as both the memory construction model and answer model, except for ablations in Table 2 which vary the answer model.
Mem0 is configured with its default settings for memory extraction and organization.
MIRIX is configured with its default multi-dimensional memory structure.
Both frameworks construct memories by processing dialogues sequentially, then answer questions by retrieving from the constructed memory bank.
D.3 Training-based Methods
Memory-R1 results are taken from the original paper. Note that Memory-R1 uses Llama-3.1-8B-Instruct as both the memory construction and answer model, which differs from our evaluation setup. For reference, our method with Qwen3-4B as both the construction and answer model achieves 82.00% on LoCoMo (Table 2), substantially outperforming Memory-R1’s reported 62.74%.
Appendix E Ablation Experiment Details
E.1 Gradient Weighting Ablation
To efficiently explore the effect of contribution-aware gradient weighting, we conduct this ablation on a reduced training set consisting of 10 dialogues sampled from the 50 RL training dialogues. We vary $\alpha\in\{1,2,4,8,16\}$ while keeping all other hyperparameters fixed. The setting $\alpha=1$ corresponds to uniform weighting without contribution-aware scaling. Results are shown in Figure 4.
E.2 Reward Density Ablation
We conduct this ablation on the full training set of 50 dialogues. To validate the importance of dense session-level rewards, we vary the reward density by randomly skipping task reward computation for a fraction of sessions. Specifically, at density $1/d$, each session independently has a probability $1/d$ of receiving a task reward. Sessions without task rewards still receive format validity and length penalty signals, but no QA-based feedback. All configurations are trained for the same number of epochs to ensure fair comparison. Results are shown in Figure 5.
Appendix F Cost Analysis
Our training pipeline incurs API costs at two stages: (1) data preparation (one-time) and (2) RL training (per-run).
F.1 Data Preparation Costs (One-Time)
Expert Trajectory Generation.
We use Claude 4.5 Sonnet to generate memory management demonstrations for 50 LongMemEval conversations (2,400 sessions). Each session invokes 4 memory agent calls (Core, Episodic, Semantic, Procedural), resulting in 9,600 API calls. Based on actual logged data from 952 agent calls, the average token consumption is 6,011 input tokens (agent prompt + existing memories + current session) and 784 output tokens (JSON memory operations). Total: $\sim$58M input + $\sim$8M output tokens. Cost: \$294.
Synthetic Question Generation.
For all 50 conversations (2,400 sessions), we generate 5 QA pairs per session using Claude 4.5 Opus. Based on analysis of generated QA files, each call uses $\sim$3,500 input tokens (QA prompt template + memory state + current session) and produces $\sim$460 output tokens (5 questions with answers in JSON). Total: $\sim$8.5M input + $\sim$1.1M output tokens. Cost: \$70.
F.2 RL Training Costs (Per-Run)
During GRPO training (2,400 sessions $\times$ 8 rollouts $\times$ 5 epochs = 96,000 session-rollouts), reward computation evaluates 5 QA questions per rollout, yielding 480,000 QA evaluations. Each evaluation involves two API calls: answer generation and LLM judging, both using GPT-4.1-mini. Answer generation consumes approximately 960M input and 48M output tokens (\$461), while LLM judging requires approximately 241M input and 15M output tokens (\$120). The per-run total is \$581.
Appendix G Case Study
To illustrate the behavioral changes discussed in Section 4.5.2, we present representative examples comparing model outputs across training stages (Base, SFT, RL) on identical inputs.
G.1 Core Memory: Precision Evolution
Scenario.
The user discusses updates to their vegetarian lifestyle. The current Core Memory contains:
Dietary preference: Exploring plant-based options
Family: Has children
Hobbies: Cooking, reading
User Message.
“I’ve been vegetarian for about 6 months now. My kids love when I make vegan pancakes on Sunday mornings. We tried oat milk last week and it’s now my favorite plant-based milk!”
Base Model.
Produces a fragmented append without structure:
{"operation": "APPEND", "content": "vegetarian 6 months. oat milk favorite. vegan pancakes."}
SFT Model.
Rewrites the entire memory block:
[…]
RL Model.
Performs a targeted replacement of only the relevant field:
{"operation": "REPLACE", "old_text": "Dietary preference: Exploring plant-based options", "new_text": "Dietary preference: Vegetarian (6 months); favorite plant milk: oat milk; makes vegan pancakes for kids on Sundays"}
Analysis.
The Base model outputs unstructured fragments. SFT learns valid formatting but defaults to REWRITE, risking information loss when reorganizing the entire block. The RL model learns that REPLACE achieves targeted updates while preserving unrelated fields (Family, Hobbies).
G.2 Episodic Memory: Temporal Chain Construction
Scenario.
The user follows up on a previously mentioned trip. The retrieved memory contains:
[EPISODIC] 2024-02-15: User planned trip to Japan | Details: Planning vacation to Japan, interested in visiting Kyoto and Tokyo. Researching cherry blossom season timing.
User Message.
“Remember when I told you about my trip to Japan last month? Well, I finally uploaded all the photos! I got some amazing shots at the bamboo forest in Kyoto.”
Base Model.
Adds a disconnected entry with no reference to the planning phase:
{"operations": [{"action": "ADD", "memory": "User went to Japan. Took photos at bamboo forest."}]}
SFT Model.
Overwrites the planning record with completion information:
{"operations": [{"action": "UPDATE", "old_memory": "2024-02-15: User planned trip to Japan...", "new_memory": "2024-03-15: User completed trip to Japan | Details: Took trip to Japan last month. Visited Kyoto and Tokyo. Took photos at bamboo forest in Arashiyama. Uploaded all photos."}]}
RL Model.
Creates a new event entry and updates the original with status annotation:
{"operations": [{"action": "ADD", "memory": "2024-03-10: User visited bamboo forest in Arashiyama, Kyoto | Details: Took amazing photos during trip to Japan; shared photos highlighting scenic beauty and personal enjoyment"}, {"action": "UPDATE", "old_memory": "2024-02-15: User planned trip to Japan...", "new_memory": "2024-02-15: User planned trip to Japan [COMPLETED] | Details: Planning vacation to Japan -> Trip completed March 2024. Visited Kyoto (bamboo forest) and Tokyo. Photos uploaded."}]}
Analysis.
The Base model creates isolated entries without temporal connections. SFT’s UPDATE operation overwrites planning details, losing information about the user’s original interests (cherry blossom season research). The RL model constructs a temporal chain: preserving the planning phase, adding the completion event as a separate entry, and marking status transitions.
G.3 Semantic Memory: Comprehensive Fact Capture
Scenario.
The user expresses interest in art after visiting a museum.
User Message.
“I’ve been really into impressionist art lately. Went to see the Monet exhibition at the museum last weekend. His water lilies series is absolutely stunning.”
Base Model.
Adds a simple entry with minimal detail:
{"operations": [{"action": "ADD", "memory": "User likes impressionist art. Saw Monet exhibition."}]}
SFT Model.
Updates an existing hobby entry:
{"operations": [{"action": "UPDATE", "old_memory": "Hobbies: Enjoys visiting museums | Details: Regular museum visitor...", "new_memory": "Hobbies: Enjoys visiting museums | Details: Regular museum visitor, interested in various exhibitions, particularly impressionist art. Recently saw Monets water lilies series exhibition at a museum last weekend."}]}
RL Model.
Creates multiple entries capturing both user preference and conceptual knowledge:
{"operations": [{"action": "ADD", "memory": "User - Art Interest: Impressionist art | Details: Recently developed strong interest in impressionist art, particularly Monets water lilies series. Describes them as absolutely stunning’."}, {"action": "ADD", "memory": "Concept - Monets Water Lilies | Details: A series of paintings by Claude Monet depicting water lilies, displayed at Musee de lOrangerie in Paris."}]}
Analysis.
The Base model captures only surface-level facts. SFT consolidates new information into existing entries, compressing the user’s art interest into a general hobby record. The RL model learns to output multiple distinct entries: one for user-specific preferences (interest in impressionist art) and another for conceptual knowledge (Monet’s Water Lilies). This separation enables more precise retrieval when answering questions about the user’s artistic taste versus factual questions about specific artworks.
G.4 Procedural Memory: Version Preservation
Scenario.
The user describes a refined coffee routine. The retrieved memory contains:
[PROCEDURAL] Users coffee brewing method | Steps: 1. Grind beans. 2. Add hot water. 3. Wait 4 minutes. 4. Press and pour. | Context: French press method.
User Message.
“I finally perfected my morning coffee routine! I grind 18g of beans, use 92°C water, bloom for 30 seconds, then do a slow pour over 3 minutes.”
Base Model.
Adds a simple entry without procedural structure:
{"operations": [{"action": "ADD", "memory": "User has new coffee method: 18g beans, 92C water, 30s bloom, 3min pour."}]}
SFT Model.
Overwrites the French press method entirely:
{"operations": [{"action": "UPDATE", "old_memory": "Users coffee brewing method | Steps: 1. Grind...", "new_memory": "Users pour-over coffee brewing method | Steps: 1. Grind 18g of beans. 2. Use 92C water. 3. Bloom for 30 seconds. 4. Slow pour over 3 minutes. | Context: Pour-over method, refined morning routine."}]}
RL Model.
Adds the new method as a versioned entry, preserving the original:
{"operations": [{"action": "ADD", "memory": "Pour-over coffee method (v2, current) | Steps: 1. Grind 18g beans to medium-fine. 2. Heat water to 92C. 3. Bloom for 30 seconds. 4. Slow pour in circular motion for 3 minutes. 5. Allow to drip. | Context: Refined morning routine, replaces previous French press method"}]}
Analysis.
This case exhibits the starkest contrast. SFT’s UPDATE completely erases the French press procedure—if the user later asks “How did I use to make coffee with my French press?”, the system cannot answer. The RL model learns through downstream QA feedback that procedural knowledge should be versioned rather than overwritten. The new entry explicitly references the previous method while the original remains retrievable.
Appendix H Additional Analysis
H.1 Validation of Retrieval-Based Attribution
To validate that retrieval frequency serves as an adequate proxy for downstream contribution, we recruited three graduate students (M.S. or above in Computer Science, familiar with dialogue systems and NLP) to independently evaluate 100 synthetic QA pairs sampled from the RL training set. For each question, annotators examined whether the dominant memory type (by retrieval count) contains information directly used to derive the correct answer, with the final label determined by majority vote. Results are shown in Table 4.
Category Count %
Dominant type contains answer-relevant information 95 95
Correct answer relies on a non-dominant type 2 2
Policy model failed to store relevant information 3 3
Table 4: Human evaluation of retrieval-based attribution on 100 sampled QA pairs.
Pairwise inter-annotator agreement ranges from 92.0% to 96.0%, with all three annotators agreeing on 92.0% of samples. While retrieval frequency is not a perfect indicator, the dominant type aligns with the actual information source in 95% of cases, confirming that it provides a probabilistically reliable attribution signal.
H.2 Synthetic QA Quality Evaluation
To assess the quality of synthetic QA pairs used for dense reward computation, we recruited three graduate students (M.S. or above in Computer Science, familiar with dialogue systems) to independently evaluate 100 randomly sampled QA pairs from the RL training set. Each pair was assessed on three criteria, with final labels determined by majority vote:
  • Question Reasonableness: whether the question is well-formed and answerable given the conversation context up to the current session.
  • Answer Correctness: whether the ground-truth answer is factually correct based on the information present in the dialogue history.
  • Answer Specificity: whether the answer is specific enough to unambiguously address the question, rather than being vague or overly general.
Results are shown in Table 5.
Criterion Pass Rate Avg. Agreement
Question Reasonableness 91% 91.3%
Answer Correctness 92% 90.7%
Answer Specificity 93% 94.0%
Table 5: Human evaluation of synthetic QA quality on 100 sampled pairs. Agreement refers to average pairwise inter-annotator agreement.
Among the 8 incorrect answers (by majority vote), the primary error types are temporal calculation errors (6/8, e.g., incorrect date arithmetic or imprecise relative time references) and over-inference across sessions (2/8, where answers combine information from multiple sessions in ways not directly supported by the dialogue).
H.3 Feasibility of Open-Source Answer Models
Our evaluation pipeline uses Claude 4.5 Sonnet as the default answer model, which incurs API costs. To assess whether open-source models can serve as viable alternatives, we fix the memory bank to that constructed by Claude 4.5 Sonnet and evaluate different answer models on LoCoMo under identical retrieval settings. This setup isolates the answer model’s capability by controlling for memory quality.
Answer Model LoCoMo
Claude 4.5 Sonnet 82.61
GPT-4.1-mini 80.88
Qwen3-30B-A3B 80.84
Table 6: Performance of different answer models on LoCoMo using memory constructed by Claude 4.5 Sonnet, isolating the effect of answer model capability.
Qwen3-30B-A3B achieves performance comparable to GPT-4.1-mini, demonstrating that open-source models of similar scale can serve as cost-effective alternatives for the answer component in both evaluation and RL reward computation.
Appendix I Prompt Templates
This section presents the prompt templates used in our framework. For readability, we have made minor formatting adjustments (e.g., line breaks and indentation) to the original prompts.
Figure 8: Prompt template for Core Memory.
Figure 9: Prompt template for Episodic Memory.
Figure 10: Prompt template for Procedural Memory.
Figure 11: Prompt template for Core Memory compression.
Figure 12: Prompt template for QA answering.
Figure 13: Prompt template for LLM judge evaluation.
Figure 14: Prompt template for synthetic question generation.